A 3.0L flask contains 3.5mol of HI, 4 Moles of H2 and 0.3 mol of I2 in equilibrium.
Calculate the Kc of the reaction.
H2(g)+I2(g) HI(g)
1
Expert's answer
2020-12-21T03:39:40-0500

"Kc = \frac{(HI)^r}{(H2)(I2)}"


"Kc = \frac{(1.6)\times(1.6)}{(1.33)(0,1)}"


= 10.11